hayhay6:

What is the equation of a line that passes through the point (5, -3) and has a slope of -2 ?

Vocaloid:

Point slope form (y-y1) = m(x-x1) Where m is the slope and (x1,y1) is your given point. Plug in the appropriate quantities. If you need y = mx + b form, solve the equation for y.